Tracy and Klein studied together for the midterm exam. It took them 4 3/4 hours from start to finish during this time they took a half hour break to eat lunch and two more breaks that were one third of an hour to get snacks what is the best estimate of the time that was spent studying?
Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

4 3/4 -1/2 for lunch=4 1/4
subtract 2(1/3)=2/3 from 4 1/4
Can make it 4 3/12-8/12, using the common denominator of 12. Borrow 5 from 12/12, and it is 3 7/12.
-----------------
Subtracting 2/3 from 4 1/4
Go the other way.
from 2/3 to 1 you need 1/3
From 1 to 4 you need 3
from 4 to 4 1/4 you need 1/4
3 +(1/3)+(1/4)=3 (4/12)+(3/12)=3 (7/12)

Answer by stanbon(75887)

Tracy and Klein studied together for the midterm exam. It took them 4 3/4 hours from start to finish during this time they took a half hour break to eat lunch and two more breaks that were one third of an hour to get snacks what is the best estimate of the time that was spent studying?
Ans:: 4 3/4 - [1/2 + 1/3 + 1/3] = 19/4 - (3+2+2)/6 = 19/4-7/6 = (114-28)/24
= 86/24 = 43/12 = 3 7/12 = 3 hrs 35 min
----------
Easier way:: 4hr 45 min - [30 + 20 + 20]min = 3hr 35 min
---------------
Cheers,
Stan H.